OpenStudy (anonymous):

substitution for y=x+6 y=2x

jimthompson5910 (jim_thompson5910):

start with y = x+6 since y = 2x, we can replace 'y' with 2x

jimthompson5910 (jim_thompson5910):

so.. y = x+6 2x = x+6 x = ???